### v2

# Pinpointing Important Genetic Variants via A Feature-level Filter and Group Knockoffs

Abstract
Identifying variants that carry substantial information on the trait of interest remains a core topic in genetic studies. In analyzing the EADB-UKBB dataset to identify genetic variants associated with Alzheimer’s disease (AD), however, we recognize that both existing marginal association tests and conditional independence tests using knockoffs suffer either power loss or lack of informativeness, especially when strong correlations exist among variants. To address the limitations of existing knockoff filters, we propose a new feature-versus-group (FVG) filter that is more powerful and precise in identifying important features from a set of strongly correlated features using group knockoffs. In extensive simulation studies, the FVG filter controls the expected proportion of false discoveries and identifies important features with enhanced power and greater precision. Applying the proposed method to the EADB-UKBB dataset, we discover important variants from 84 loci (same as the most powerful group knockoff filter) with catching sets of substantially smaller size and higher purity.
Keywords: Alzheimer’s disease genetics, False discovery rate, Feature selection, Knockoffs.
1 Introduction
1.1 Conditional Independence Test to Discover Disease-associated Variants
Testing conditional independence between a set of features $\textbf{X}=(X_{1},\ldots,X_{p})^{\mathsf{T}}$ and the response of interest $Y$ is an important topic in various research areas, including causal inference (Peters,, 2015; Cai et al.,, 2022), genetic analysis (Khera and Kathiresan,, 2017; Zhu et al.,, 2018) and graphical model learning (Deka et al.,, 2016; Tugnait,, 2022). In the era of big data, as the number of features increases, the need for statistical approaches for simultaneous inference of conditional independence between hundreds of thousands of features and the response keeps emerging.
Specifically, developing statistical methods for conditional independence tests with type-I error rate control has been an important and popular research topic of large-scale genome-wide association studies (GWAS; Tang and He,, 2021; Hou et al.,, 2023; Morra et al.,, 2023) in recent years. With the ultimate goal of identifying novel targets for the development of genomic-driven medicine, the contribution of more genetic variants to complex phenotypes has been investigated in GWAS. However, our ability to detect causal genetic variants and to translate discoveries into insights of the underlying genetic mechanisms does not increase proportionally to the growth of the scale of GWAS. On one hand, with the stringent criterion to control the family-wise error rate (FWER), conventional GWAS are usually suboptimal in terms of the statistical power of detecting important variants. On the other hand, because conventional GWAS base their inference on marginal association models which regress $Y$ on one $X_{j}$ at a time, they can only provide blurred results with a lot of proxy variants correlated with the true causal variants (Schaid et al.,, 2018). Both issues become more severe when more variants are sequenced at a higher resolution because the $p$-value threshold of FWER control also becomes more stringent and there are more proxy variants strongly correlated with true causal variants. Although many fine-mapping methods have been proposed (Benner et al.,, 2017; Zou et al.,, 2022) to pinpoint the causal variants, they can only partially address the aforementioned issues without control of type-I error rate, calling for new statistical inference methods with better theoretical properties and empirical performance in multiple testing of conditional independence for real-world genetic data analysis.
In this article, we investigate the stage I meta-analysis of the European Alzheimer & Dementia Biobank (EADB) dataset and the UK Biobank (UKBB) dataset (Bellenguez et al.,, 2022) to identify genetic variants associated with AD. This meta-analyzed EADB–UKBB dataset is deposited in the European Bioinformatics Institute GWAS Catalog [^1] https://www.ebi.ac.uk/gwas/ (Accession number: GCST90027158), including $39,106$ clinically diagnosed AD cases, $46,828$ proxy cases, and $401,577$ control cases from $15$ European countries. We focus on $640,576$ directly genotyped variants with minor allele frequency (MAF) at least $0.01$, resulting in a processed dataset $\mathbb{G}$ of $487,511$ observations and $640,576$ features. With such a large dataset, we aim to discover as many important genetic variants that carry independent information of AD without making too many false discoveries. Mathematically, such a task can be transferred into a multiple testing problem of conditional independence between a large set of variants $\textbf{X}=(X_{1},\ldots,X_{p})^{\mathsf{T}}$ and the binary response $Y$ which indicates whether a case is clinically diagnosed AD.
1.2 Existing Approaches and Their Limitations
Recent decades have witnessed fruitful works in the development of multiple testing procedures. To control the family-wise error rate (FWER) of making at least one false discovery, a lot of $p$-value-based methods have been developed as improvements of the classic Bonferroni correction, including the Šidák correction (Šidák,, 1967), Holm’s step-down procedure (Holm,, 1979) and Hochberg’s step-up procedure (Hochberg,, 1988). However, FWER has been criticized for its conservativeness in many settings, especially when most of the signals are weak. To perform powerful inference, Benjamini and Hochberg, (1995) proposed the false discovery rate (the expected proportion of false discoveries in all discoveries) as an alternative type-I error rate measure and developed a procedure to perform multiple testing with provable FDR control under the assumption that $p$-values are independent or positively dependent. Inspired by this pioneering work, Benjamini and Yekutieli, (2001) developed a generalized procedure without assumptions on dependencies among $p$-values, while Storey, (2002) and Whittemore, (2007) investigated the Bayesian counterpart of FDR. However, as most of the aforementioned methods require $p$-values of nulls for all features, their feasibility on multiple testing of conditional independence becomes questionable when the number of features $p$ greatly exceeds the sample size $n$. The main reason is that $p$-values of conditional independence under high-dimensional scenarios are nontrivial and generally not available except in a very few specific cases (Tibshirani et al.,, 2016; Lee et al.,, 2016).
Unlike all the above methods, the recently developed knockoff filter (Barber and Candès,, 2015; Candès et al.,, 2018) and its variants (Gimenez et al.,, 2019; Gimenez and Zou,, 2019; Barber and Candès,, 2019; Huang and Janson,, 2020; Ren and Candès,, 2023) are a family of approaches that can control finite-sample FDR in multiple testing of conditional independence hypotheses at the feature-versus-feature level
$H^{(\text{ff})}_{j}:X_{j}\perp Y|\textbf{X}_{-j},\text{ where }\textbf{X}_{-j}=(X_{1},\ldots,X_{j-1},X_{j+1},\ldots,X_{p})^{\mathsf{T}},\quad j=1,\ldots,p,$ (1)
without any assumption on the conditional distribution $Y|\textbf{X}$. Specifically, these methods first construct knockoffs $\widetilde{\textbf{X}}$ that mimic the dependency among features X and are not dependent to the response $Y$. By doing so, knockoffs play the role of synthetic controls: if $H^{(\text{ff})}_{j}$ is true, ${X}_{j}$ and $\widetilde{X}_{j}$ are indistinguishable even with information of $Y$; if $H^{(\text{ff})}_{j}$ is false, ${X}_{j}$ and $\widetilde{X}_{j}$ possess different dependencies with $Y$. By comparing dependencies between $Y$ and X with dependencies between $Y$ and $\widetilde{\textbf{X}}$, these methods have both controlled FDR and superior power in detecting true causal features as validated in extensive simulation studies.
However, in the analysis of real-world genetic data, methods that infer $H^{(\text{ff})}_{j}$’s usually suffer severe power loss. The main reason is that linkage disequilibrium could introduce high correlations among genetic variants, making it difficult to distinguish variants that are truly associated with the response (whose $H^{(\text{ff})}_{j}$’s are false) from their strongly correlated null variants (whose $H^{(\text{ff})}_{j}$’s are true). This would significantly decrease the number of discoveries as shown by the simulation studies in Barber and Candès, (2015) and Candès et al., (2018). Such a power loss issue would deteriorate when more genetic variants with strong correlations are recorded under a growing resolution of sequencing. For example, when we apply the model-X knockoff filter (Candès et al.,, 2018) to the EADB-UKBB dataset (Bellenguez et al.,, 2022), only 138 AD-associated genetic variants from 49 loci (here, we define two loci as different if they are at least 1Mb away from each other) are identified under the target FDR level $0.10$ as shown in Table 1. Compared with the 51 loci discovered by marginal association test ($p$-value threshold: $5\times 10^{-8}$) in Bellenguez et al., (2022), although model-X knockoff filter identified 6 new loci, it misses several important AD loci with strong signals, including the “FAM193B” locus, the “FAM157C” locus, the “NSF” locus, the “SIGLEC11” locus and the “APP” locus.
Table 1: Summary of results by applying different methods to the EADB-UKBB dataset.
Method Number of Average number of Average Average
identified variants per size of purity of
loci identified locus catching sets catching sets
Marginal association test 51 18.078 18.078 0.477
Model-X knockoff filter 49 2.816 1.000 1.000
Group knockoff filter 84 23.714 13.105 0.610
: Catching sets refer to
(a) sets of variants within different loci identified by marginal association test;
(b) variants identified by the model-X knockoff filter;
(c) variant groups identified by the group knockoff filter.
: Purity refers to the minimum absolute correlation within the catching set.
To address this power loss issue, Dai and Barber, (2016) proposed the group knockoff filter. By defining groups $B_{1},\ldots,B_{k}$ such that features with strong correlations are in the same group and changing the inference target as
$H^{\text{(gg)}}_{k}:\textbf{X}_{B_{k}}\perp Y|\textbf{X}_{-B_{k}},\quad k=1,\ldots,K,$ (2)
at the group-versus-group level, important features and their strongly correlated null features can be rejected as a group. Because the grouping is done to ensure features from different groups are weakly correlated, the group knockoff filter can circumvent the obstacle of high correlation in inferring $H^{(\text{ff})}_{j}$ with significantly higher power. To apply the group knockoff filter to the EADB-UKBB dataset (Bellenguez et al.,, 2022), we first compute correlation $\text{cor}(G_{i},G_{j})$ between any pair of variants $G_{i}$ and $G_{j}$ and construct $321,535$ variant groups by applying the hierarchical clustering (average linkage with cutoff value $0.5$) on the distance matrix $(1-|\text{cor}(G_{i},G_{j})|)_{p\times p}$. By doing so, variants in the EADB-UKBB dataset (Bellenguez et al.,, 2022) are grouped such that the average absolute correlation of variants from different groups is lower than 0.5. Under such a group structure, the group knockoff filter identifies 152 AD-associated variant groups from 84 loci under the target FDR level $0.10$. Specifically, among these 152 identified groups, 126 are located outside the APOE/APOC region with the strongest association to AD (chromosome 19, positions 44909011 $\sim$ 45912650), including groups in loci “FAM193B”, “FAM157C”, “NSF”, “SIGLEC11” and “APP” that are missed by the model-X knockoff filter. In addition, the group knockoff filter also identifies 39 new loci that are not discovered by Bellenguez et al., (2022).
However, doing so can only provide a series of AD-associated variant groups without information on which variants contribute more to AD variation within each group, leaving the target of identifying AD-associated variants not fully addressed. This can be observed from Table 1. With an average size $13.105$ of the identified groups given by the group knockoff filter, 35 identified groups contain at least 10 genetic variants and the 5 largest groups contain 314, 232, 136, 121 and 117 variants respectively. With such large sizes, which genetic variants in these groups contribute more to AD variation remains unanswered.
1.3 Our Contribution
In this article, to address the limitations of existing knockoff filters in discovering important variants of AD in the analysis of the EADB-UKBB dataset, we propose a new knockoff filter that is more powerful in inferring conditional independence hypotheses and identifying important features from a set of strongly correlated features. Our contribution in this article is two-folded.
  • Presenting a family of conditional independence hypotheses that allow identifying important features using group knockoffs. Compared to existing conditional independence hypotheses, the new hypotheses allow the use of group knockoffs construction mechanism (Dai and Barber,, 2016) to provide informative discoveries at the level of features with higher power.
  • Developing the new knockoff filter with FDR control, high empirical power and informativeness [^2] In this article, informativeness refers to the precision of catching sets in identifying conditionally independent signals and is quantified by the size and purity of catching sets.. Specifically, we thoroughly investigate the theoretical properties of the developed knockoff filter and empirically evaluate their performance via simulated experiments.
The rest of this article is organized as follows. In Section 2, we develop our knockoff filter and perform theoretical analysis. Section 3 illustrates the empirical performance of our method and its advantage over existing alternatives in power and informativeness via simulated experiments. Finally, we apply our method to the AD analysis of the EADB-UKBB dataset described in Section 4 and demonstrate that the method exhibits high power and precision. Section 5 concludes with discussions.
2 Methodology
2.1 Problem Statement
Consider independent and identically distributed (i.i.d.) samples $\{(\textbf{x}_{i},y_{i})|i=1,\ldots,n\}$ from a joint distribution $f(\mathbf{X},Y)$, where features $\textbf{X}=(X_{1},\ldots,X_{p})^{\mathsf{T}}$ are partitioned into $K$ disjoint groups $B_{1},\ldots,B_{K}$. Let $\mathbb{X}=[\textbf{x}_{1},\ldots,\textbf{x}_{n}]^{\mathsf{T}}$ denote the data matrix and $\textbf{y}=(y_{1},\ldots,y_{n})^{\mathsf{T}}$ be the vector of responses. Our target is to perform multiple testing on conditional independence hypotheses,
$H_{j}^{(\text{fg})}:X_{j}\perp Y|\textbf{X}_{-B_{k}},\quad k=1,\ldots,K,\quad j\in B_{k},$ (3)
where $\textbf{X}_{-B_{k}}=(\textbf{X}^{\mathsf{T}}_{B_{1}},\ldots,\textbf{X}^{\mathsf{T}}_{B_{k-1}},\textbf{X}^{\mathsf{T}}_{B_{k+1}},\ldots,\textbf{X}^{\mathsf{T}}_{B_{K}})^{\mathsf{T}}$ is the subvector of X created by excluding all features in the same group $B_{k}$ of $X_{j}$. Here, we refer $H_{j}^{(\text{fg})}$ as the conditional independence hypothesis at the feature-versus-group level as it depicts the independence between a feature and the response conditional on other feature groups which omits information of other features in the same group. Under the belief that the response $Y$ only depends on a relatively small number of features, we would like to find as many important features without making too many false discoveries with respect to $H_{j}^{(\text{fg})}$’s. In other words, our target is to obtain a rejection set $\mathcal{R}^{(\text{fg})}=\{j|H_{j}^{(\text{fg})}\text{ is rejected}\}$ such that the false discovery rate (FDR, the expectation of false discovery proportion),
$\text{FDR}^{(\text{fg})}=\mathbf{E}\{\text{FDP}^{(\text{fg})}\},\quad\text{where }\text{FDP}^{(\text{fg})}=\frac{\#(\mathcal{R}^{(\text{fg})}\cap\mathcal{H}_{0}^{(\text{fg})})}{1\vee\#\mathcal{R}^{(\text{fg})}},\quad\mathcal{H}_{0}^{(\text{fg})}=\{j|H_{j}^{(\text{fg})}\text{ is true}\},$
is controlled under the target level $\alpha$ ($10<\alpha<$ 1).
Under the same setup of Candès et al., (2018) that the distribution of features X is completely known while no information is provided regarding the conditional distribution $Y|\textbf{X}$, our method differs in the hypotheses to test (3) from existing approaches, which test either conditional independence hypotheses (1) at the feature-versus-feature level or conditional independence hypotheses (2) at the group-versus-group level. Such a difference is the origin of our method’s advantages over the existing knockoff filters. On one hand, unlike the group knockoff filter (Dai and Barber,, 2016), the conclusion provided by our method is at the feature level and is thus more informative on which features contribute more to the variation of $Y$. On the other hand, compared to those that infer $H^{(\text{ff})}_{j}$’s, our method can also utilize group knockoffs that are more liberal in construction to gain power in selecting important features. However, such advantages do not come free as new challenge is posted that no existing filters can provide control on $\text{FDR}^{(\text{fg})}$, which is elaborated and addressed in the following sections.
2.2 Feature Importance Scores Under Group Knockoffs
To infer conditional independence hypotheses (3) with group knockoffs, our inference procedure, similar to the existing ones, include four main steps as follows.
  • $\diamond$
    (Knockoffs Construction): Construct the knockoff copy $\widetilde{\textbf{X}}$ of X such that $\perp Y|\textbf{X}$ X and the joint distribution of $(\textbf{X},\widetilde{\textbf{X}})$ is exchangeable at the group level (Dai and Barber,, 2016),
    $(\textbf{X},\widetilde{\textbf{X}})_{\text{swap}(\mathcal{S})}{\displaystyle\mathop{=\joinrel=}^{D}}(\textbf{X},\widetilde{\textbf{X}}),\quad\forall\mathcal{S}\subset\{1,\ldots,K\},$ (4)
    where $(\textbf{X},\widetilde{\textbf{X}})_{\text{swap}(\mathcal{S})}$ is obtained by swapping $\textbf{X}_{B_{k}}$ and $\widetilde{\textbf{X}}_{B_{k}}$ in $(\textbf{X},\widetilde{\textbf{X}})$ for all $k\in\mathcal{S}$.
  • $\diamond$
    (Feature Importance Scores Calculation): For each $H^{(\text{fg})}_{j}$, compute feature importance scores $T_{j}$ for feature $X_{j}$ and $\widetilde{T}_{j}$ for its knockoff copy $\widetilde{X}_{j}$.
  • $\diamond$
    (Feature Statistics Calculation): For each $H^{(\text{fg})}_{j}$, summarize importance scores $T_{j}$ and $\widetilde{T}_{j}$ as a feature statistic $W_{j}=w_{j}(T_{j},\widetilde{T}_{j})$ using an anti-symmetric function $w_{j}$ such that $w_{j}(T_{j},\widetilde{T}_{j})=-w_{j}(\widetilde{T}_{j},T_{j})$.
  • $\diamond$
    (Feature Filtering): Reject $H^{(\text{fg})}_{j}$ if $W_{j}$ is positive and large enough (i.e., $T_{j}$ dominates $\widetilde{T}_{j}$).
However, directly inputting $W_{1},\ldots,W_{p}$ to the model-X knockoff filter at feature level (Candès et al.,, 2018) cannot provide control on $\text{FDR}^{(\text{fg})}$ of the rejection set. The main reason is that under the group knockoffs model $f(\textbf{X},\widetilde{\textbf{X}})$, $W_{1},\ldots,W_{p}$ do not possess the coin-flipping property (i.e., conditional on $|W_{j}|$’s and $\text{sign}(W_{j})$’s for those false $H^{(\text{fg})}_{j}$’s, $\text{sign}(W_{j})$’s for those true $H^{(\text{fg})}_{j}$’s are independently and uniformly distributed on {±$\{\pm 1\}$), the key property that guarantees FDR control in Candès et al., (2018). Specifically, for most of the widely-used feature importance scores (e.g., the marginal correlations, the Lasso coefficient-difference statistics and the Lasso signed max statistics; Barber and Candès,, 2015, Candès et al.,, 2018), $W_{j}$’s for those true $H^{(\text{fg})}_{j}$’s are dependent coin flips. This comes from (4) where exchangeability only exists among groups, making signs of $W_{j}$’s from the same group not independent.
Thus, we relax the unachievable coin-flipping property to the between-group coin-flipping property which can be satisfied by feature importance scores $W_{j}$’s in the form of
$\begin{cases}T_{j}=t([\mathbb{X}_{j},\mathbb{X}_{B_{k}\setminus\{j\}},\widetilde{\mathbb{X}}_{j},\widetilde{\mathbb{X}}_{B_{k}\setminus\{j\}},\{\mathbb{X}_{-B_{k}},\widetilde{\mathbb{X}}_{-B_{k}}\}],\textbf{y}),\\ \widetilde{T}_{j}=t([\widetilde{\mathbb{X}}_{j},\widetilde{\mathbb{X}}_{B_{k}\setminus\{j\}},{\mathbb{X}}_{j},{\mathbb{X}}_{B_{k}\setminus\{j\}},\{\mathbb{X}_{-B_{k}},\widetilde{\mathbb{X}}_{-B_{k}}\}],\textbf{y}),\end{cases}\quad k=1,\ldots,K\text{, }j\in B_{k},$ (5)
where $\text{where }\{\mathbb{X}_{-B_{k}},\widetilde{\mathbb{X}}_{-B_{k}}\}\text{ denotes the set of unordered pairs }\{\mathbb{X}_{j^{\prime}},\widetilde{\mathbb{X}}_{j^{\prime}}\}\text{ for all }j^{\prime}\notin B_{k},$
as shown in Theorem 1.
Theorem 1.
For any feature importance scores $T_{j}$ and $\widetilde{T}_{j}$ in the form of (5), feature statistics $W_{j}=w_{j}(T_{j},\widetilde{T}_{j})$ ($j=1,\ldots,p$) with antisymmetric functions $w_{j}$’s satisfy the between-group coin-flipping property that
  1. $\star$
    conditional on $|W_{1}|,\ldots,|W_{p}|$ and $\text{\rm sign}(W_{j})$’s for those false $H^{(\text{fg})}_{j}$’s,
    1. $\diamond$
      (Uniformity) $\text{\rm sign}(W_{j})$ is uniformly distributed on {±$\{\pm 1\}$ for all $j\in\mathcal{H}_{0}^{(\text{fg})}$;
    2. $\diamond$
      (Between-group independence) for any $k\neq k^{\dagger}$, $\text{\rm sign}(W_{j})$ and $\text{\rm sign}(W_{j^{\dagger}})$ are independent for any $j\in B_{k}\cap\mathcal{H}_{0}^{(\text{fg})}$ and $j^{\dagger}\in B_{k^{\dagger}}\cap\mathcal{H}_{0}^{(\text{fg})}$,
if $H_{j}^{(\text{fg})}$ implies $H_{k}^{\text{(gg)}}$ for all $j\in B_{k}$ and $k=1,\ldots,K$.
The proof of Theorem 1 is provided in Section A of the Appendix. Specifically, Theorem 1 relies on a condition “$H_{j}^{(\text{fg})}$ implies $H_{k}^{\text{(gg)}}$ for all $j\in B_{k}$ and $k=1,\ldots,K$”. Such a condition is generally valid in genetic data analysis, analogous to the usual expectation in fine-mapping that the association between an important variant $X_{j}$ and the response $Y$ is in the same direction as the direct effect of $X_{j}$ on $Y$.
Examples of feature importance scores in the form of (5) include:
  • $\diamond$
    marginal correlation with response: absolute values of the empirical marginal correlations $T_{j}=\Big{|}\widehat{\rho}(X_{j},Y)\Big{|}$ and $\widetilde{T}_{j}=\Big{|}\widehat{\rho}(\widetilde{X}_{j},Y)\Big{|}$;
  • $\diamond$
    joint lasso: absolute values of lasso estimators $T_{j}=\Big{|}\widehat{\beta}_{j}\Big{|}$ and $\widetilde{T}_{j}=\Big{|}\widehat{\widetilde{\beta}}_{j}\Big{|}$ of the linear model
    $Y=\sum_{j=1}^{p}\left({\beta}_{j}{{X}}_{j}+\widetilde{\beta}_{j}\widetilde{{X}}_{j}\right)+e;$ (6)
  • $\diamond$
    marginal correlation with lasso residual: absolute values of estimated marginal correlations $T_{j}=\Big{|}\widehat{\rho}(X_{j},\hat{e})\Big{|}$ and $\widetilde{T}_{j}=\Big{|}\widehat{\rho}(\widetilde{X}_{j},\hat{e})\Big{|}$ between $(X_{j},\widetilde{X}_{j})^{\mathsf{T}}$ and the lasso residual of the linear model
    $Y=\sum_{j^{\dagger}\notin B_{k}}\left({\beta}_{j^{\dagger}}{{X}}_{j^{\dagger}}+\widetilde{\beta}_{j^{\dagger}}\widetilde{{X}}_{j^{\dagger}}\right)+e,\quad\text{where }j\in B_{k};$
  • $\diamond$
    separate lasso: absolute values of lasso estimators $T_{j}=\Big{|}\widehat{\beta}_{j}\Big{|}$ and $\widetilde{T}_{j}=\Big{|}\widehat{\widetilde{\beta}}_{j}\Big{|}$ of the separate linear model
    $Y={\beta}_{j}{{X}}_{j}+\widetilde{\beta}_{j}\widetilde{{X}}_{j}+\sum_{j^{\dagger}\notin B_{k}}\left({\beta}_{j^{\dagger}}{{X}}_{j^{\dagger}}+\widetilde{\beta}_{j^{\dagger}}\widetilde{{X}}_{j^{\dagger}}\right)+e,\quad\text{where }j\in B_{k}.$
Although the above feature importance scores can be used to compute feature statistics $W_{1},\ldots,W_{j}$ with the between-group coin-flipping property, they suffer several problems in practice. On one hand, with not much information of $H_{j}^{(\text{fg})}$’s, marginal correlation with response usually lacks power. So do the joint lasso feature importance scores, which do not directly depict evidence against $H_{j}^{(\text{fg})}$’s. On the other hand, with the need to perform $p$ and $K$ lasso linear model fittings, separate lasso and marginal correlation with lasso residual are both computationally intensive.
Faced with such a tradeoff between computational efficiency and power, we propose the combined feature importance scores as computationally efficient proxies of the separate lasso importance scores
$\begin{cases}T_{j}=|g_{j}|\cdot|\widehat{\rho}(X_{j},Y)|,\\ \widetilde{T}_{j}=|g_{j}|\cdot|\widehat{\rho}(\widetilde{{X}}_{j},Y)|,\\ \end{cases}\quad k=1,\ldots,K\text{, }j\in B_{k},$ (7)
where $g_{j}$ is computed as follows.
  1. 1.
    Fit the joint linear model (6) and compute $\widehat{\gamma}_{j}=(\widehat{\beta}_{j}+\widehat{\widetilde{\beta}}_{j})/2$ and $\widehat{\delta}_{j}=(\widehat{\beta}_{j}-\widehat{\widetilde{\beta}}_{j})/2$ as the approximate estimator of the reparametrized model
    $Y=\sum_{j=1}^{p}{\gamma}_{j}\left({{X}}_{j}+\widetilde{{X}}_{j}\right)+\sum_{j=1}^{p}{\delta}_{j}\left({{X}}_{j}-\widetilde{{X}}_{j}\right)+e.$
  2. 2.
    For $k=1,\ldots,K$ and $j\in B_{k}$, compute
    $g_{j}$ $=\widehat{\gamma}_{j}+\sum_{j^{\dagger}\in B_{k}\setminus\{j\}}\frac{\psi_{j,j^{\dagger}}-\boldsymbol{\Psi}_{j,-B_{k}}\boldsymbol{\Psi}^{-1}_{-B_{k},-B_{k}}\boldsymbol{\Psi}_{-B_{k},j^{\dagger}}}{\psi_{j,j}-\boldsymbol{\Psi}_{j,-B_{k}}\boldsymbol{\Psi}^{-1}_{-B_{k},-B_{k}}\boldsymbol{\Psi}_{-B_{k},j}}\widehat{\gamma}_{j^{\dagger}},$
    where $\boldsymbol{\Psi}=(\psi_{j,j^{\dagger}})_{p\times p}=\boldsymbol{\Sigma}-\textbf{D}/2,\quad\boldsymbol{\Sigma}=\text{Var}(\textbf{X}),\quad\textbf{D}=\boldsymbol{\Sigma}-\text{Cov}(\textbf{X},\widetilde{\textbf{X}}),$
    as the approximate estimator of the model
    $Y={g}_{j}\left({{X}}_{j}+\widetilde{{X}}_{j}\right)+\sum_{j^{\prime}\notin B_{k}}{\gamma}_{j^{\prime}}\left({{X}}_{j^{\prime}}+\widetilde{{X}}_{j^{\prime}}\right)+{h}_{j}\left({{X}}_{j}-\widetilde{{X}}_{j}\right)+\sum_{j^{\prime}\notin B_{k}}{\delta}_{j^{\prime}}\left({{X}}_{j^{\prime}}-\widetilde{{X}}_{j^{\prime}}\right)+e.$
    In other words, ${g}_{j}$ measures the overall contribution of ${X}_{j}$ and $\widetilde{{X}}_{j}$ to the variation of $Y$ beyond features from other groups.
By doing so, feature statistics $W_{j}=w_{j}(T_{j},\widetilde{T}_{j})$ ($j=1,\ldots,p$) with antisymmetric functions $w_{j}$’s satisfy the between-group coin-flipping property because
$\text{sign}(W_{j})=\text{sign}(w_{j}(|\widehat{\rho}({{X}}_{j},Y)|,|\widehat{\rho}(\widetilde{{X}}_{j},Y)|)),$
whose between-group coin-flipping property is validated in Theorem 1. In addition, as $g_{j}$ is derived to measure the total dependency between $(X_{j},\widetilde{{X}}_{j})$ and the response $Y$ conditional on $(\textbf{X}_{-B_{k}},\widetilde{\textbf{X}}_{-B_{k}})$, $|W_{j}|$’s computed from the combined feature importance scores provide proxies of $|W_{j}|$’s computed from the separate lasso feature importance scores.
2.3 Naive Feature Filters under Group Knockoffs
By Theorem 1, feature statistics $W_{j}$’s only satisfy the between-group coin-flipping property and thus directly applying the model-X knockoff filter (Candès et al.,, 2018) would violate the FDR control. However, we can still utilize the between-group independence among $W_{j}$’s for inference with FDR control as follows.
Recognizing that independence remains true among $W_{j}$’s for those true $H^{(\text{fg})}_{j}$’s from different groups, we consider aligning $W_{j}$’s in Table 2 such that $W_{j}$’s for all features in the group $B_{k}$ are in the $k$-th column and $|W_{(k1)}|\geq|W_{(k2)}|\geq\cdots$. Let $\mathcal{C}_{l}$ be the set of features whose feature statistics are aligned in the $l$-th row and consider the rejection set in the form of $\mathcal{R}^{(\text{fg})}(t)=\{j|W_{j}\geq t\}$.
Table 2: Alignment of feature statistics $W_{j}$’s such that different groups correspond to different columns. Here, $W_{(kl)}$ corresponds to the feature from the $k$-th group that is aligned in the $l$-th row.
Row $B_{1}$ $B_{2}$ $B_{3}$cdots
1 $W_{(11)}$ $W_{(21)}$ $W_{(31)}$cdots
[…]
  • On one hand, we find that within each row, there exists at most one $W_{j}$ from each group. As feature statistics $W_{j}$’s satisfy the between-group coin-flipping property, we have that $W_{j}$’s within the same row satisfy the coin-flipping property introduced in Candès et al., (2018). Thus, we can estimate the false discovery proportion (FDP) of the rejection subset $\mathcal{R}^{(l)}(t)=\{j\in\mathcal{C}_{l}|W_{j}\geq t\}$ as
    $\widehat{\text{FDP}}^{(l)}(t)=\frac{1+\#\{j\in\mathcal{C}_{l}|W_{j}\leq-t\}}{1\vee\#\{j\in\mathcal{C}_{l}|W_{j}\geq t\}},$ (8)
    for any $t>0$ and $l=1,2\ldots$ in the analogous way of Candès et al., (2018).
  • On the other hand, we have $\text{FDP}^{(\text{fg})}(t)$ of $\mathcal{R}^{(\text{fg})}(t)$ is a weighted mean of $\text{FDP}^{(l)}(t)$’s of subsets $\mathcal{R}^{(l)}(t)=\mathcal{R}^{(\text{fg})}(t)\cap\mathcal{C}_{l}$ ($l=1,2,\ldots$) that
    $\text{FDP}^{(\text{fg})}(t)=$$\frac{\#(\mathcal{R}^{(\text{fg})}(t)\cap\mathcal{H}_{0}^{(\text{fg})})}{1\vee\#\mathcal{R}^{(\text{fg})}(t)}$
    $=$ $\sum_{l=1}^{\phi(t)}\frac{\#(\mathcal{R}^{(l)}(t)\cap\mathcal{H}_{0}^{(\text{fg})})}{1\vee\#\mathcal{R}^{(\text{fg})}(t)}$
    $=$ $\sum_{l=1}^{\phi(t)}\frac{1\vee\#\mathcal{R}^{(l)}(t)}{1\vee\#\mathcal{R}^{(\text{fg})}(t)}\times\frac{\#(\mathcal{R}^{(l)}(t)\cap\mathcal{H}_{0}^{(\text{fg})})}{1\vee\#\mathcal{R}^{(l)}(t)}$
    $=$ $\sum_{l=1}^{\phi(t)}\frac{1\vee\#\mathcal{R}^{(l)}(t)}{1\vee\#\mathcal{R}^{(\text{fg})}(t)}\times\text{FDP}^{(l)}(t),$
    where the second equality comes from the fact that rejection subsets $\mathcal{R}^{(l)}(t)$ can be nonempty only for the first $\phi(t)=\max_{k}\#\{j\in B_{k}||W_{j}|\geq t\}$ rows because $|W_{(k1)}|\geq|W_{(k2)}|\geq\cdots$ with each group $B_{k}$.
As a result, using $\widehat{\text{FDP}}^{(l)}(t)$’s in (8), we can estimate $\text{FDP}^{(\text{fg})}(t)$ of $\mathcal{R}^{(\text{fg})}(t)$ by
$\widehat{\text{FDP}}^{(\text{fg})}(t)$ $=\sum_{l=1}^{\phi(t)}\frac{1\vee\#\mathcal{R}^{(l)}(t)}{1\vee\#\mathcal{R}^{(\text{fg})}(t)}\times\widehat{\text{FDP}}^{(l)}(t)$
$=\sum_{l=1}^{\phi(t)}\frac{1\vee\#\{j\in\mathcal{C}_{l}|W_{j}\geq t\}}{1\vee\#\{j|W_{j}\geq t\}}\times\frac{1+\#\{j\in\mathcal{C}_{l}|W_{j}\leq-t\}}{1\vee\#\{j\in\mathcal{C}_{l}|W_{j}\geq t\}}$
$=\frac{{\phi(t)}+\#\{j|W_{j}\leq-t\}}{1\vee\#\{j|W_{j}\geq t\}},$
leading to the naive filter (Algorithm 1).
Algorithm 1 Naive feature filter with group knockoffs.
1:  Input: Groups $B_{1},\ldots,B_{K}$, feature statistics $W_{1},\ldots,W_{p}$ and the target level $\alpha>0$.
2:  Compute
$t_{\alpha}=\min\Biggl{\{}t>0\Bigg{|}\widehat{\text{FDP}}^{(\text{fg})}(t)=\frac{{\phi(t)}+\#\{j|W_{j}\leq-t\}}{1\vee\#\{j|W_{j}\geq t\}}\leq\alpha\Biggr{\}},$ (9)
where $\phi(t)=\max_{k}\#\{j\in B_{k}||W_{j}|\geq t\}$.
3:  Output: The rejection set $\mathcal{R}^{(\text{fg})}=\{j|W_{j}\geq t_{\alpha}\}$.
2.4 Rigorous Feature Filters under Group Knockoffs
Although the naive filter seems intuitive in FDR control, such a control can not be proven. To fill the gap, we provide an alternative interpretation of Algorithm 1 and consider a variation that provides provable FDR control as follows.
With the target to reject as many false $H_{j}^{(\text{fg})}$’s as possible while $\text{FDR}^{(\text{fg})}$ is under control, multiple testing of $H_{j}^{(\text{fg})}$’s can be transferred into an optimization problem. Based on the construction of feature statistics that $W_{j}$’s for those false $H_{j}^{(\text{fg})}$’s tend to be positive and large, Algorithm 1 solves the following optimization problem
[…]
$\mathcal{R}^{(\text{fg})}=\cup_{l}\mathcal{R}^{(l)}=\cup_{l}\{j\in\mathcal{C}_{l}|W_{j}\geq t^{(l)}\},\quad t^{(1)}=t^{(2)}=\cdots.$
Based on such an observation, we utilize theoretical results of Katsevich and Sabatti, (2019) that for any $t>0$,
$\textbf{E}\left[\sup_{t>0}\frac{\#\{j\in\mathcal{H}_{0}^{(\text{fg})}\cap\mathcal{C}_{l}|W_{j}\geq t\}}{1+\#\{j\in\mathcal{H}_{0}^{(\text{fg})}\cap\mathcal{C}_{l}|W_{j}\leq-t\}}\right]\leq 1.93,$
to modify the above optimization problem as
[…]
$\mathcal{R}^{(\text{fg})}=\cup_{l}\mathcal{R}^{(l)}=\cup_{l}\{j\in\mathcal{C}_{l}|W_{j}\geq t^{(l)}\},$
where $v_{l}$’s are pre-calculated budgets for contribution of $\mathcal{R}^{(l)}$’s to $\widehat{\text{FDP}}^{(\text{fg})}$. By doing so, the rejection set $\mathcal{R}^{(\text{fg})}$ obtained by solving (10) has controlled $\text{FDR}^{(\text{fg})}$ as shown in Theorem 2.
Theorem 2.
The rejection set $\mathcal{R}^{(\text{fg})}$ obtained by solving (10) controls $\text{FDR}^{(\text{fg})}$ at the target level $\alpha>0$.
The proof of Theorem 2 is provided in Section B of the Appendix. Similar to the discussion in Sesia et al., (2020), the correction factor $1.93$ is required in the proof for technical reasons, while empirical evidence suggests that it may be practically unnecessary. To solve the optimization problem (10), we develop the feature-versus-group (FVG) filter (Algorithm 2) in an analogous manner to the GGM knockoff filter (Li and Maathuis,, 2021).
We also provide other variations for FDR control with details in Section C of the Appendix.
Algorithm 2 Feature-versus-group (FVG) filter with group knockoffs.
1:  Input: Groups $B_{1},\ldots,B_{K}$, feature statistics $W_{1},\ldots,W_{p}$ and the target level $\alpha>0$.
2:  Align $W_{j}$’s as shown in Table 2 such that $W_{j}$’s for all features in the group $B_{k}$ are in the $k$-th column and $|W_{(k1)}|\geq|W_{(k2)}|\geq\cdots$.
3:  Compute budget $v_{l}$ for the $l$-th row using $|W_{1}|,\ldots,|W_{p}|$ for $l=1,2,\ldots$ such that $\sum_{l=1}^{\infty}v_{l}=1$.
4:  Compute grids $\mathcal{G}_{l}=\{1/v_{l},2/v_{l},\ldots,({neg}_{l}+1)/v_{l}\}$ for the $l$-th row where $neg_{l}=\#\{j\in\mathcal{C}_{l}|W_{j}<0\}$.
5:  Combine $\mathcal{G}_{comb}=(\cup_{l}\mathcal{G}_{l})\cup\{0\}$ and sort values ${grid}_{(1)}>{grid}_{(2)}>\ldots$ in $\mathcal{G}_{comb}$.
6:  Initialize $b=0$
7:  repeat
8:     Update $b=b+1$.
9:     Compute
$t^{(l)}=\min\Biggl{\{}t>0\Bigg{|}\frac{1+\#\{j\in\mathcal{C}_{l}|W_{j}\leq-t\}}{v_{l}}\leq{grid}_{(b)}\Biggr{\}},$ (11)
for $l=1,2,\ldots$.
10:  until $\text{I}\left(\max_{j\in\mathcal{C}_{l}}|W_{j}|\geq t\right)\times\frac{1+\#\{j\in\mathcal{C}_{l}|W_{j}\leq-t^{(l)}\}}{1\vee[\sum_{l}\#\{j\in\mathcal{C}_{l}|W_{j}\geq t^{(l)}\}]}\leq\frac{v_{l}\alpha}{1.93}$ for all $l$.
11:  Output: The rejection set $\mathcal{R}^{(\text{fg})}=\cup_{l}\{j\in\mathcal{C}_{l}|W_{j}\geq t^{(l)}\}$.
Remark 1.
In the FVG filter (Algorithm 2), the budgets $v_{l}$’s can be computed adaptively to the prior belief of the number of important features within each groups. Here, we provide two strategies of specifying $v_{l}$’s.
  • If it is believed that all important groups have similar effect size on the response $Y$, we suggest letting
    $v_{l}=\frac{\sum_{k}|W_{(kl)}|}{\sum_{l}\sum_{k}|W_{(kl)}|},\quad l=1,2,\ldots,$
    such that contribution of $\mathcal{R}^{(l)}$’s to $\widehat{\text{FDP}}^{(\text{fg})}$ are constrained by the sum of signal magnitudes $\sum_{k}|W_{(kl)}|$ in the $l$-th row of Table 2.
  • If it is believed that most of important groups have small effect size while a small number of important groups have large effect size, we suggest letting
    $v_{l}=\frac{\sum_{k}|W_{(kl)}|/l}{\sum_{l}\sum_{k}|W_{(kl)}|/l},\quad l=1,2,\ldots,$
    such that more budgets are allocated to the first several rows of Table 2 and more important features in important groups with small effect sizes are identified.
3 Simulated Experiments
To evaluate the proposed FVG filter in both FDR control and identifying important features, we conduct extensive experiments under various scenarios. Specifically, throughout simulated experiments, we generate group knockoffs under the maximum entropy (ME) construction of Chu et al., (2023), compute feature statistics as $W_{j}=\text{sign}\left(T_{j}-\widetilde{T}_{j}\right)\cdot\max\left(T_{j},\widetilde{T}_{j}\right)$ with the joint lasso importance scores.
3.1 Simulated Experiments with Synthetic Data
To examine how the FVG filter performs under different sample sizes, we simulate 1000 datasets with a fixed number of features $p=250$ for sample sizes $n=500,1000$ and $2000$. Specifically, we partition these $250$ features into $K=50$ groups of equal size $B_{1}=\{X_{1},\ldots,X_{5}\},\ldots,B_{50}=\{X_{246},\ldots,X_{250}\}$. Based on such a partition, we generate for each simulated dataset $n$ feature vectors $\textbf{x}_{1},\ldots,\textbf{x}_{n}$ from the multivariate normal distribution,
$\text{MVN}\left\{\mathbf{0},\begin{pmatrix}0.3\cdot\textbf{I}+0.7\cdot\mathbf{1}\mathbf{1}^{\mathsf{T}}&0.3\cdot\mathbf{1}\mathbf{1}^{\mathsf{T}}&\cdots&0.3\cdot\mathbf{1}\mathbf{1}^{\mathsf{T}}\\ 0.3\cdot\mathbf{1}\mathbf{1}^{\mathsf{T}}&0.3\cdot\textbf{I}+0.7\cdot\mathbf{1}\mathbf{1}^{\mathsf{T}}&\cdots&0.3\cdot\mathbf{1}\mathbf{1}^{\mathsf{T}}\\ \vdots&\vdots&\ddots&\vdots\\ 0.3\cdot\mathbf{1}\mathbf{1}^{\mathsf{T}}&0.3\cdot\mathbf{1}\mathbf{1}^{\mathsf{T}}&\cdots&0.3\cdot\textbf{I}+0.7\cdot\mathbf{1}\mathbf{1}^{\mathsf{T}}\\ \end{pmatrix}\right\}.$ (12)
Given features $\textbf{x}_{1},\ldots,\textbf{x}_{n}$, responses $y_{1},\ldots,y_{n}$ are simulated from the linear model
$Y=\sum_{k=1}^{K}\textbf{X}_{B_{k}}^{\mathsf{T}}\boldsymbol{\beta}_{B_{k}}+e,\quad e\sim\text{N}(0,4^{2}),$ (13)
where
$\boldsymbol{\beta}_{B_{k}}=\begin{cases}(\textbf{1}_{2}^{\mathsf{T}},-0.2\cdot\textbf{1}_{3}^{\mathsf{T}})^{\mathsf{T}},&k=1,2,\\ (2.5,-1.2\cdot\textbf{1}_{2}^{\mathsf{T}},\textbf{0}_{2}^{\mathsf{T}})^{\mathsf{T}},&k=3,4,\\ (0.3\cdot\textbf{1}_{2}^{\mathsf{T}},\textbf{0}_{3}^{\mathsf{T}})^{\mathsf{T}},&k=5,\ldots,10\\ \textbf{0}_{5},&k=11,\ldots,50,\\ \end{cases}$ (14)
Under (12)-(14), we have $H_{j}^{(\text{fg})}$’s are false for $j=51,\ldots,250$ and there are 28 important features. Empirical FDR and power of the FVG filter in identifying important features under different sample sizes are visualized in Figure 1. For comparison, we also evaluate the performance of model-X knockoff filter (Candès et al.,, 2018), group knockoff filter (Dai and Barber,, 2016) and the multilayer knockoff filter (Katsevich and Sabatti,, 2019) in FDR control and identifying important features.
(a) $n=500$.
(b) $n=1000$.
(c) $n=2000$.
Figure 1: Empirical FDR and power of FVG filter with respect to the target FDR level ($\alpha$) over 1000 simulated datasets of different sample sizes in comparison to model-X knockoff filter (Candès et al.,, 2018), group knockoff filter (Dai and Barber,, 2016) and the multilayer knockoff filter (Katsevich and Sabatti,, 2019).
It is clear that the FVG filter manages to control $\text{FDR}^{(\text{fg})}$ under any target level $\alpha$. For a fixed target level $\alpha$, we find the empirical power increases to $1$ as $n$ grows, suggesting that the FVG filter is consistent in identifying important features. Although existing methods can also control FDR and have growing power, they suffer different types of performance deterioration in identifying important features. On one hand, group knockoff filter suffers power loss when the target FDR level is lower than $0.10$ due to the threshold issue mentioned in Gimenez and Zou, (2019). Even with power improvement as the target FDR level grows from $0.10$ and $0.20$, the informativeness of rejection sets obtained by group knockoff filter is poor as all important groups ($B_{1},\ldots,B_{10}$) are rejected as a whole without distinguishing which features are more important. As a comparison, the average size of nonempty catching sets $\mathcal{R}_{k}=B_{k}\cap\mathcal{R}$ ($k=1,\ldots,50$) obtained by FVG filter is less than $2$ under the target level $\alpha=0.1$. On the other hand, although the multilayer knockoff filter has power improvement over model-X knockoff filter by focusing important feature selection on those identified feature groups, there remains a significant power gap between FVG filter and the multilayer knockoff filter when the target FDR is large.
3.2 Simulated Experiments with Real-world Genetic Data
To mimic the dependency structure among features in real-world genetic analysis and investigate how the proposed method performs in real applications, we simulate 500 datasets based on $6095$ variants in the APOE/APOC region (chromosome 19, positions 44909011 $\sim$ 45912650) from the whole-genome sequencing data (NG00067.v5) of the Alzheimer’s Disease Sequencing Project (Leung et al.,, 2019). Specifically, we restrict our study to $6952$ individuals with estimated European ancestry rate [^3] Estimated by SNPWeights v2.1 (Chen et al.,, 2013) using reference populations from the 1000 Genomes Consortium (The 1000 Genomes Project Consortium,, 2015). greater than 80% and $p=1000$ variants $X_{1},\ldots,X_{p}$ whose minor allele frequency (MAF) is larger than $1\%$ and pairwise correlations are in $[-0.95,0.95]$ to avoid the existence of statistically indistinguishable variants with nearly-perfect correlations. Among these variants, we compute the correlation $\text{cor}(X_{i},X_{j})$ between any pair of variants $(X_{i},X_{j})$ over all 6952 individuals and construct variant groups ${B}_{1},\ldots,{B}_{K}$ by applying the hierarchical clustering (average linkage with cutoff value $0.5$) on the distance matrix $(1-|\text{cor}(X_{i},X_{j})|)_{p\times p}$. By randomly sampling $n=5000$ individuals without replacement, each simulated dataset is obtained by collecting the sampled individuals’ variants and generating responses $y_{1},\ldots,y_{n}$ from the linear model
$Y=\sum_{j=1}^{p}X_{j}{\beta}_{j}+e,\quad e\sim\text{N}(0,4),$ (15)
where only $k=10$ randomly selected coefficients ${\beta}_{j}$’s are nonzero and follow $N(0,0.225)$. As a result, important features refer to variants $X_{j}$’s whose corresponding coefficients $\beta_{j}$’s are nonzero and important groups are those containing at least one important feature. Specifically, most important groups only contain one important feature.
Figure 2: Empirical performance of the model-X knockoff filter (Candès et al.,, 2018), the group knockoff filter (Dai and Barber,, 2016), the multilayer knockoff filter (Katsevich and Sabatti,, 2019) and our FVG filter over 500 simulated data: (a) empirical FDR and power with respect to the target FDR level; (b) average size of catching sets; (c) average purity of catching sets.
To illustrate the empirical benefit of the FVG filter, we apply it to find important features in these 500 simulated datasets with comparison to the model-X knockoff filter (Candès et al.,, 2018), the group knockoff filter (Dai and Barber,, 2016) and the multilayer knockoff filter (Katsevich and Sabatti,, 2019). Empirical FDR and power of all methods are presented in Figure 2 (a). For fair comparison, we use the joint lasso feature importance score when implementing both existing methods and our algorithm. It is found that our FVG filter can empirically control FDR. Compared to the model-X knockoff filter which can only distinguish around $50\%$ of important features when the target FDR level is relaxed to $20\%$, our FVG filter, the multilayer knockoff filter and the group knockoff filter are able to identify about $70\%\sim 80\%$ of important features (or feature groups) when the target FDR level is $10\%$. However, the inference results given by the multilayer knockoff filter and the group knockoff filter are far from being informative and precise. To illustrate this, we compute catching set $\mathcal{R}_{k}$ of each variant group from the rejection set $\mathcal{R}$, where
  • $\mathcal{R}_{k}=B_{k}$ if $H^{\text{(gg)}}_{k}$ is rejected by the group knockoff filter;
  • $\mathcal{R}_{k}=B_{k}\cap\mathcal{R}$ if the rejection set $\mathcal{R}$ is obtained from the model-X knockoff filter or our FVG filter.
  • For each group $B_{k}$, we let
    • $\mathcal{R}_{k}=B_{k}$ if $H^{\text{(gg)}}_{k}$ but all $H^{\text{(ff)}}_{j}$’s ($j\in B_{k}$) is rejected by the multilayer knockoff filter;
    • $\mathcal{R}_{k}=B_{k}\cap\mathcal{R}$ if $H^{\text{(gg)}}_{k}$ and some $H^{\text{(ff)}}_{j}$’s ($j\in B_{k}$) are rejected by the multilayer knockoff filter.
We then calculate the average size and purity (minimum absolute correlation) of catching sets to measure the informativeness and precision of $\mathcal{R}$ given by different filters in Figure 2 (b)-(c). Here, it is desired that sizes of catching sets are smaller as more unimportant variants are excluded. Even when there remain multiple features in catching sets, high purity is desirable to exclude unimportant variants weakly correlated to the important ones (or poor-quality proxy variants). In contrast to the group knockoff filter whose average size and purity of catching sets are almost $6$ and smaller than $20\%$ respectively (which means that catching sets obtained by the group knockoff filter contain a lot of poor-quality proxy variants), the precision of our FVG filter in pinpointing true important variants is close to the model-X knockoff filter and is higher than the multilayer knockoff filter. With an average size close to $1$ and average purity greater than $80\%$, our FVG filter can directly obtain informative catching sets containing only one or two features in most cases. In summary, our FVG filter manage to simultaneously achieve high power and high informativeness in identifying important features.
4 AD Analysis of EADB–UKBB Dataset
Given the promising performance of the proposed filter in simulated experiments, we apply our method to the EADB-UKBB dataset where model-X knockoff filter exhibits power loss and the group knockoff filter returns not informative enough results in Section 1.2.
To infer which variants are associated with AD, we first compute correlation $\text{cor}(G_{i},G_{j})$ between any pair of variants $G_{i}$ and $G_{j}$ and construct variant groups $B_{1},\ldots,B_{K}$ by applying the hierarchical clustering (average linkage with cutoff value $0.5$) on the distance matrix $(1-|\text{cor}(G_{i},G_{j})|)_{p\times p}$. By doing so, we obtain the same $321,535$ groups $B_{1},\ldots,B_{k}$ as the group knockoff filter in Section 1.2 where the average absolute correlation of variants from different groups are lower than 0.5. Based on the correlation matrix $(\text{cor}(G_{i},G_{j}))_{p\times p}$ and variant groups $B_{1},\ldots,B_{K}$, we generate group knockoffs $\widetilde{\mathbb{G}}$ under the maximum entropy (ME) construction of Chu et al., (2023). Given the response of interest $\textbf{y}=(y_{1},\ldots,y_{n})^{\mathsf{T}}$ where $y_{i}=1$ if the $i$-th observation corresponds to a clinically diagnosed AD case and $y_{i}=0$ otherwise, we use absolute values of lasso estimators of the logistic regression model,
$\text{logit}\{\text{Pr}(Y=1|{\textbf{G}},\widetilde{\textbf{G}})\}=\beta_{0}+\sum_{j=1}^{p}\left\{{G}_{j}{\beta}_{j}+\widetilde{G}_{j}\widetilde{\beta}_{j}\right\}$
as feature importance scores. In other words, we compute $T_{j}=\left|\widehat{\beta}_{j}\right|$ and $\widetilde{T}_{j}=\left|\widehat{\widetilde{\beta}}_{j}\right|$, which are also used in other knockoff-based methods and provide more informative rejection sets without large power loss given the large sample size. We then calculate feature statistics $W_{j}=\text{sign}\left(T_{j}-\widetilde{T}_{j}\right)\cdot\max\left(T_{j},\widetilde{T}_{j}\right)$ and implement the FVG filter with target FDR level $\alpha=0.10$. Variants in the rejection set $\mathcal{R}^{(\text{fg})}$ under target FDR level $\alpha=0.10$ are presented in Figure 3 with their post-selection importance scores $W_{j}\cdot\text{I}(W_{j}>0)$ and names of the closest genes.
Figure 3: Manhattan plot of genetic variants under our FVG filter. Identified variants under target FDR level $\alpha=0.10$ are highlighted with names of their closest genes.
Details of variants identified by the FVG filter under the target FDR level $0.10$ are provided in Table 3. In total, we identify 205 variants that contribute to AD variation substantially, which comes from 154 groups in 84 loci. Similar to the literature, multiple variants are identified in the APOE/APOC region with the strongest association to AD (chromosome 19, positions 44909011 $\sim$ 45912650). In addition, we also manage to identify variants rs744373 and rs2118506 close to gene “BIN1”, variant rs1532278 close to gene “CLU”, variant rs10792832 close to gene “PICALM”, variants rs4277405, rs4309 and rs3730025 close to gene “ACE”, all of which are also reported in He et al., (2022) and Bellenguez et al., (2022). Compared with the model-X knockoff filter, the FVG filter only misses 2 loci identified by the model-X knockoff filter but discover 37 more loci, which suggests that our method would not suffer from the same power loss issue of the model-X knockoff filter.
Table 3: Details of variants identified by the FVG filter under the EADB–UKBB dataset and target FDR level $\alpha=0.10$.
Identified Position Closest Identified Position Closest Identified Position Closest Identified Position Closest
Variant Gene Variant Gene Variant Gene Variant Gene
Chromosome 1 rs3997700 41251889 TREML5P rs9516168 93085490 LINC00363 rs3852860 44879709 NECTIN2
rs11579514 161173522 B4GALT3 rs9296558 47484147 CD2AP Chromosome 14 rs365653 44858389 NECTIN2
rs782791 203029691 PPFIA4 rs10948363 47520026 CD2AP rs74093831 105733666 ELK2AP rs73052307 44881148 NECTIN2
rs4562624 207512620 CR1 rs12525381 696327 PTPRK rs873533 106667442 LINC00221 rs11668327 44895376 TOMM40
rs3818361 207611623 CR1 rs117591349 78478451 MIR10524 rs17125924 52924962 FERMT2 rs449647 44905307 APOE
rs6701713 207612944 CR1 rs313161 85332752 LINC02535 rs4904929 92470949 SLC24A4 rs10420036 44882772 NECTIN2
rs4844610 207629207 CR1 Chromosome 7 Chromosome 15 rs283813 44885917 NECTIN2
Chromosome 2 rs221834 100745552 ZAN rs59685680 50709337 SPPL2A rs7254892 44886339 NECTIN2
rs116038905 105805908 NCK2 rs144867634 111940111 DOCK4 rs28470808 50735410 SPPL2A rs7412 44908822 APOE
rs74851408 117201103 DDX18 rs10952168 1554792 TMEM184A rs2043085 58388755 LIPC rs445925 44912383 APOC1
rs6733839 127135234 BIN1 rs56402156 143406388 TCAF2 rs593742 58753575 ADAM10 rs390082 44913574 APOC1
rs730482 127136908 BIN1 rs12703526 143410495 TCAF2 rs117618017 63277703 APH1B rs190712692 44921921 APOC1
rs744373 127137039 BIN1 rs9640386 143435090 TCAF2 rs3816385 64126456 SNX1 rs41290108 44874585 NECTIN2
rs745717 127136321 BIN1 rs2189965 28132395 JAZF1 rs56117790 79075361 RASGRF1 rs73033507 44928146 APOC1P1
rs2118506 127139927 BIN1 rs2122623 37842715 NME8 Chromosome 16 rs79701229 44881674 NECTIN2
rs10496731 134840058 ACMSD rs7809394 55077689 EGFR rs12325539 30022312 C16orf92 rs157580 44892009 TOMM40
rs4954187 134842848 ACMSD rs12669393 7815748 UMAD1 rs11865499 31120929 KAT8 rs2075649 44892073 TOMM40
rs3819121 134865290 ACMSD rs4725076 8204779 ICA1 rs8058370 81738205 CMIP rs8106922 44898409 TOMM40
rs6725887 202881162 WDR12 rs1476679 100406823 ZCWPW1 rs12446759 81739398 CMIP rs405697 44901434 TOMM40
rs10933431 233117202 INPP5D rs41280968 100105324 AP4M1 rs11548656 81883307 PLCG2 rs7259620 44904531 TOMM40
rs7421448 233117495 INPP5D rs13228694 100342684 STAG3L5P-PVRIG2P-PILRB rs1071644 81937798 PLCG2 rs440446 44905910 APOE
rs35349669 233159830 INPP5D Chromosome 8 rs56407236 90103687 FAM157C rs6857 44888997 NECTIN2
rs3771371 71400409 ZNF638 rs2436889 102564430 ODF1 Chromosome 17 rs2075650 44892362 TOMM40
rs113068192 76786226 LRRTM4 rs4731 11808828 FDFT1 rs79759095 44451394 GPATCH8 rs157581 44892457 TOMM40
Chromosome 3 rs79832570 144042819 SPATC1 rs5910 44372421 ITGA2B rs34404554 44892652 TOMM40
rs7634492 136105288 PPP2R3A rs61732533 144053248 OPLAH rs199439 46716137 NSF rs75627662 44910319 APOE
rs16824536 155069722 MME rs34674752 144099319 SHARPIN rs199451 46724418 NSF rs438811 44913484 APOC1
rs61762319 155084189 MME rs73223431 27362470 PTK2B rs2671659 49397167 PHB rs283814 44885967 NECTIN2
rs4974180 56200107 ERC2 rs1532278 27608798 CLU rs616338 49219935 ABI3 rs439401 44911194 APOE
Chromosome 4 rs9297949 94957217 NDUFAF6 rs2072153 49312652 ZNF652 rs59325138 44913034 APOC1
rs6448453 11024404 CLNK Chromosome 9 rs2526378 58326988 TSPOAP1-AS1 rs584007 44913221 APOC1
rs13116567 12239152 LINC02270 rs1131773 93077974 SUSD3 rs2632516 58331728 TSPOAP1-AS1 rs157595 44922203 APOC1
rs502746 14019593 LINC01182 Chromosome 10 rs9896864 63458947 CYB561 rs4803770 44924096 APOC1P1
rs116791081 143428212 GAB1 rs7920721 11678309 ECHDC3 rs4277405 63471557 ACE rs71352239 44926286 APOC1P1
Chromosome 5 rs7068614 11679478 ECHDC3 rs4309 63482562 ACE rs118060185 44928417 APOC1P1
rs25992 14707491 ANKH rs10764258 20634336 MIR4675 rs3730025 63480412 ACE rs144311893 44920687 APOC1
rs871269 151052827 TNIP1 rs117980033 29966853 JCAD rs9901675 7581494 CD68 rs114533385 44933496 APOC1P1
rs335473 177559423 FAM193B rs1227759 69831131 COL13A1 Chromosome 19 rs79429216 44942260 APOC4-APOC2
rs4700716 180138792 RASGEF1C rs6586028 80494228 TSPAN14 rs445752 43956053 ZNF221 rs12721109 44943964 APOC4-APOC2
rs10068419 86923485 LINC02059 Chromosome 11 rs2571174 44430314 ZNF229 rs16979595 44974124 CLPTM1
Chromosome 6 rs74685827 121482368 SORL1 rs2293164 44475997 ZNF180 rs10411314 45224801 EXOC3L2
rs976271 114361563 HDAC2-AS2 rs2298813 121522975 SORL1 rs1625502 44540845 CEACAM22P rs75987583 45745299 MEIOSIN
rs2130357 27919052 OR2B2 rs3781837 121578263 SORL1 rs10405071 44732632 BCL3 rs875121 49948251 SIGLEC11
rs13408 28244970 ZKSCAN4 rs3740890 130385218 ZBTB44-DT rs2965169 44747899 BCL3 rs1710398 51223655 CD33
rs3130931 31167111 POU5F1 rs17439299 15599040 LINC02751 rs116480598 44753996 BCL3 rs3865444 51224706 CD33
rs3132451 31614248 AIF1 rs1685404 47222114 DDB2 rs28399654 44813331 BCAM rs12974390 54264111 LILRB5
rs532965 32610196 HLA-DRB1 rs2293579 47419207 PSMC3 rs10419669 44794812 CBLC rs10404195 5908958 VMAC
rs9469112 32447376 HLA-DRA rs1582763 60254475 MS4A4E rs10405693 44823407 BCAM rs4147918 1058177 RDH13
rs9271162 32609938 HLA-DRB1 rs11824773 60309467 MS4A4A rs73572039 44839373 NECTIN2 19:1051137_CTG_C 1051138 RDH13
rs9273349 32658092 HLA-DQB1 rs17854357 65834089 SNX32 rs77301115 44893716 TOMM40 rs3752246 1056493 RDH13
rs9275184 32686937 HLA-DQB1 rs536841 86076782 PICALM rs112849259 44894050 TOMM40 rs35917007 1849148 REXO1
rs9275614 32716480 HLA-DQA2 rs10792832 86156833 PICALM rs79398853 44895528 TOMM40 Chromosome 20
rs114812713 41066261 OARD1 rs3844143 86139201 PICALM rs75687619 44896087 TOMM40 rs6064392 56409712 CASS4
rs7748513 41160234 TREM2 Chromosome 12 rs76366838 44896639 TOMM40 rs6069749 56446519 CASS4
rs2093395 41187288 TREML2 rs17860282 51346566 CELA1 rs115881343 44899959 TOMM40 Chromosome 21
rs143332484 41161469 TREM2 Chromosome 13 rs3865427 44877704 NECTIN2 rs62222965 26225443 APP
rs9394778 41247320 TREML5P
Figure 4: Comparison of our FVG filter and the group knockoff filter in analyzing the EADB–UKBB dataset: (a) Venn diagram of variant groups identified by different filters; (b) bar plot of size of catching sets identified by different filters (columns); and (c) histogram of purity of catching sets identified by different filters.
Compared to the group knockoff filter which identifies 152 AD-associated variant groups from 84 loci under the same target FDR level $0.10$, our FVG filter does not suffer power loss. As shown in Figure 4 (a), among all variant groups identified by either the group knockoff filter or our FVG filter [^4] Here, we say a variant group is identified by our FVG filter if it contains at least one variant identified by our FVG filter ., 143 groups are identified by both methods while 11 groups and 9 groups are only identified by our filter and the group knockoff filter respectively. This suggests the consistency of inference results between our FVG filter and the group knockoff filter. However, our filter manages to detect signals at variant rs1532278 close to gene “CLU”, while none of variant groups in the “CLU” locus are identified by the group knockoff filter. Without significant power loss, our method is shown to succeed in pinpointing AD-associated variants with higher precision. As shown in Figure 4 (b)-(c), all catching sets obtained by our method are of size no larger than $6$ while most of catching sets (129/154) are of size 1 and purity greater than $90\%$. In other words, most of catching sets only contain the important variants or strongly correlated proxy variants. In contrast, 37 catching sets obtained by the group knockoff filter (Dai and Barber,, 2016) contain more than $8$ variants and are far from being informative with strong impurity. Specifically, for the five largest catching sets obtained by the group knockoff filter shown in Table 4, our FVG filter manages to distinguish one or two variants from more than 100 proxy variants in each. This suggests the ability of the proposed method in pinpointing AD-associated variants. Comprehensive comparisons of our method with all existing methods are presented in Table 5.
Table 4: List of variants identified by the FVG filter in the five largest catching sets obtained by the group knockoff filter.
Identified Chromosome Position Catching sets obtained by Closest
Variant the group knockoff filter (size) Gene
rs13408 6 28244970 6:28017819:28917607:9 (121) ZKSCAN4
rs3130931 6 31167111 6:30798168:31571217:20 (136) POU5F1
rs3132451 6 31614248 6:31571218:32682663:7 (232) AIF1
rs9469112 6 32447376 6:31571218:32682663:76 (171) HLA-DRA
rs199439 17 46716137 17:43056905:45876021:36 (314) NSF
rs199451 17 46724418
Table 5: Summary of results by applying different methods on the EADB-UKBB dataset.
Method Number of Average number of Average Average
identified variants per size of purity of
loci identified locus catching sets catching sets
Marginal association test 51 18.078 18.078 0.477
Model-X knockoff filter 49 2.816 1.000 1.000
Group knockoff filter 84 23.714 13.105 0.610
FVG filter 84 2.440 1.257 0.958
5 Discussions
As more genetic variants are sequenced using the rapidly developing whole-genome sequencing technology, it is now possible to explain more diseases’ heritability via large-scale genetic studies. Noting that existing knockoff methods either lose power or lack informativeness because of strong correlations among variants in Alzheimer’s disease analysis of the EADB-UKBB dataset, we developed a new filter to select important features using group knockoffs. Specifically, we first define a new family of conditional independence hypotheses that allow inference at feature level under group knockoffs construction. Based on theoretical properties of test statistics when group knockoffs are used, we develop the FVG knockoff filter with FDR control. Extensive simulated experiments (with both synthetic data and real-world genetic data) empirically validate the FDR control of our proposed filter. Compared with the existing knockoff filters, our FVG filter is shown to be powerful in identifying important features with higher precision. When applying the proposed method to AD analysis of the real-world EADB–UKBB dataset, we identify as many signals as the group knockoff filter with as high precision as the model-$X$ knockoff filter at the same target FDR level $0.10$. Within several variant groups identified by the group knockoff filter, our method precisely distinguishes the variants that contribute more to AD variation.
Simultaneous inference of conditional independence at different layers of groups is of great necessity as we need to balance the power and the resolution (informativeness) of groups (Katsevich and Sabatti,, 2019; Sesia et al.,, 2020; Gablenz and Sabatti,, 2023). For example, in genetic analysis, variant groups with higher resolution are typically smaller and have higher correlations, making it challenging to identify important signals, but easier to interpret if they are identified. Thus, it is interesting to incorporate the proposed filter into the multilayer testing framework (Barber and Ramdas,, 2017; Katsevich and Sabatti,, 2019; Sesia et al.,, 2020; Gablenz and Sabatti,, 2023) to perform simultaneous inference of $H^{(\text{fg})}$’s at multiple layers of group structure. In addition, it is also of great interest to adopt the proposed filter in causal inference. By elaborately designing group knockoffs that can eliminate confounding effects, the proposed filter can perform causal features selection with provable FDR control when unmeasured confounders exist. We leave these tasks as future works.
Appendix A Proof of Theorem 1
Without loss of generality, we let $k=1$, $k^{\dagger}=2$, $j\in B_{1}$ and $j^{\dagger}\in B_{2}$.
  • $\star$
    (Uniformity):
    Since $H_{j}^{(\text{fg})}$ implies $H_{k}^{\text{(gg)}}$ where $j\in B_{k}$, if $j\in\mathcal{H}_{0}^{(\text{fg})}$, we have $H_{1}^{\text{(gg)}}$ is true (and $H_{j^{\prime}}^{(\text{fg})}$ is true for all $j^{\prime}\in B_{k}$) and $(X_{j},\textbf{X}_{B_{1}\setminus\{j\}})\perp Y|\textbf{X}_{-B_{1}}$ by (2). Because $(\widetilde{X}_{j},\widetilde{\textbf{X}}_{B_{1}\setminus\{j\}},\widetilde{\textbf{X}}_{-B_{1}})\perp Y|(X_{j},\textbf{X}_{B_{1}\setminus\{j\}},\textbf{X}_{-B_{1}})$, we have
    $(X_{j},\textbf{X}_{B_{1}\setminus\{j\}},\widetilde{X}_{j},\widetilde{\textbf{X}}_{B_{1}\setminus\{j\}},\widetilde{\textbf{X}}_{-B_{1}})\perp Y|\textbf{X}_{-B_{1}},$
    and thus
    $(X_{j},\textbf{X}_{B_{1}\setminus\{j\}},\widetilde{X}_{j},\widetilde{\textbf{X}}_{B_{1}\setminus\{j\}})\perp Y|(\textbf{X}_{-B_{1}},\widetilde{\textbf{X}}_{-B_{1}}).$ (16)
    By (4), we have conditional on $(\textbf{X}_{-B_{1}},\widetilde{\textbf{X}}_{-B_{1}})$,
    $(X_{j},\textbf{X}_{B_{1}\setminus\{j\}},\widetilde{X}_{j},\widetilde{\textbf{X}}_{B_{1}\setminus\{j\}}){\displaystyle\mathop{=\joinrel=}^{D}}(\widetilde{X}_{j},\widetilde{\textbf{X}}_{B_{1}\setminus\{j\}},X_{j},\textbf{X}_{B_{1}\setminus\{j\}}).$ (17)
    Thus, (16)-(17) lead to
    $([\mathbb{X}_{j},\mathbb{X}_{B_{1}\setminus\{j\}},\widetilde{\mathbb{X}}_{j},\widetilde{\mathbb{X}}_{B_{1}\setminus\{j\}},\{\mathbb{X}_{-B_{1}},\widetilde{\mathbb{X}}_{-B_{1}}\}],\textbf{y})$ (18)
[…]
    Note that swapping $\mathbb{X}_{j},\mathbb{X}_{B_{1}\setminus\{j\}}$ and $\widetilde{\mathbb{X}}_{j},\widetilde{\mathbb{X}}_{B_{1}\setminus\{j\}}$ does not change $|W_{1}|,\ldots,|W_{p}|$ and $\text{\rm sign}(W_{j})$ for those false $H^{(\text{fg})}_{j}$’s, (5) and (18) implies that conditional on $|W_{1}|,\ldots,|W_{p}|$ and $\text{\rm sign}(W_{j})$ for those false $H^{(\text{fg})}_{j}$’s,
    $(T_{j},\widetilde{T}_{j}){\displaystyle\mathop{=\joinrel=}^{D}}(\widetilde{T}_{j},T_{j}),$ (19)
[…]
  • $\star$
    (Between-group Independence):
    By (4), we have
    $({\textbf{X}}_{B_{1}},\widetilde{\textbf{X}}_{B_{1}},{\textbf{X}}_{B_{2}},\widetilde{\textbf{X}}_{B_{2}})|(\textbf{X}_{-(B_{1}\cup B_{2})},\widetilde{\textbf{X}}_{-(B_{1}\cup B_{2})})$
[…]
    If $j,j^{\dagger}\in\mathcal{H}_{0}^{(\text{fg})}$, by (5) and (19), we have conditional on $|W_{1}|,\ldots,|W_{p}|$ and $\text{\rm sign}(W_{j})$ for those false $H^{(\text{fg})}_{j}$’s,
    $({T}_{j},\widetilde{T}_{j},{T}_{j^{\dagger}},\widetilde{T}_{j^{\dagger}}){\displaystyle\mathop{=\joinrel=}^{D}}({T}_{j},\widetilde{T}_{j},\widetilde{T}_{j^{\dagger}},{T}_{j^{\dagger}}){\displaystyle\mathop{=\joinrel=}^{D}}(\widetilde{T}_{j},{T}_{j},{T}_{j^{\dagger}},\widetilde{T}_{j^{\dagger}}){\displaystyle\mathop{=\joinrel=}^{D}}(\widetilde{T}_{j},{T}_{j},\widetilde{T}_{j^{\dagger}},{T}_{j^{\dagger}}).$
    and thus $\text{\rm sign}(W_{j})$ and $\text{\rm sign}(W_{j^{\dagger}})$ are independent conditional on $|W_{1}|,\ldots,|W_{p}|$ and $\text{\rm sign}(W_{j})$ for those false $H^{(\text{fg})}_{j}$’s.
Appendix B Proof of Theorem 2
As feature statistics $W_{j}$’s satisfy the between-group coin-flipping property, we have that $W_{j}$’s within each row of Table 2 satisfy the coin-flipping property introduced in Candès et al., (2018). In other words, for each $l$, $\text{sign}(W_{j})$’s for those $j\in\mathcal{C}_{l}\cap\mathcal{H}_{0}^{(\text{fg})}$ are i.i.d. coin flips conditional on $\{W_{j}|j\in\mathcal{C}_{l}\}$ and $\{\text{\rm sign}(W_{j})|j\in\mathcal{C}_{l}\setminus\mathcal{H}_{0}^{(\text{fg})}\}$. Thus, by Lemma 3 of Katsevich and Sabatti, (2019), we have
$\textbf{E}\left\{\sup_{t>0}\frac{V^{(l)}_{\mathcal{H}_{0},+}(t)}{1+V^{(l)}_{\mathcal{H}_{0},-}(t)}\right\}\leq 1.93,\quad\text{where }\begin{cases}V^{(l)}_{\mathcal{H}_{0},+}(t)=\#\{j\in\mathcal{C}_{l}\cap\mathcal{H}_{0}^{(\text{fg})}|W_{j}\geq t\},\\ V^{(l)}_{\mathcal{H}_{0},-}(t)=\#\{j\in\mathcal{C}_{l}\cap\mathcal{H}_{0}^{(\text{fg})}|W_{j}\leq-t\}.\\ \end{cases}$ (20)
Thus, the FDR of the rejection set $\mathcal{R}^{(\text{fg})}$ obtained by solving (10) satisfies
${\text{FDR}}^{(\text{fg})}=$ $\mathbf{E}\Biggl{\{}\frac{\#(\mathcal{R}^{(\text{fg})}\cap\mathcal{H}_{0}^{(\text{fg})})}{1\vee\#\mathcal{R}^{(\text{fg})}}\Biggr{\}}$
$=$ $\mathbf{E}\Biggl{\{}\frac{\sum_{l=1}^{\infty}\text{I}\left(\max_{j\in\mathcal{C}_{l}}|W_{j}|\geq t\right)\cdot V^{(l)}_{\mathcal{H}_{0},+}(t^{(l)})}{1\vee(\sum_{l}\{j\in\mathcal{C}_{l}|W_{j}\geq t^{(l)}\})}\Biggr{\}}$
$=$ $\mathbf{E}\Biggl{\{}\sum_{l=1}^{\infty}\text{I}\left(\max_{j\in\mathcal{C}_{l}}|W_{j}|\geq t\right)\cdot\frac{1+V^{(l)}_{\mathcal{H}_{0},-}(t^{(l)})}{1\vee(\sum_{l}\{j\in\mathcal{C}_{l}|W_{j}\geq t^{(l)}\})}\cdot\frac{V^{(l)}_{\mathcal{H}_{0},+}(t^{(l)})}{1+V^{(l)}_{\mathcal{H}_{0},-}(t^{(l)})}\Biggr{\}}$
$\leq$ $\mathbf{E}\Biggl{\{}\sum_{l=1}^{\infty}\text{I}\left(\max_{j\in\mathcal{C}_{l}}|W_{j}|\geq t\right)\cdot\frac{1+\#\{j\in\mathcal{C}_{l}|W_{j}\leq-t^{(l)}\}}{1\vee[\sum_{l}\#\{j\in\mathcal{C}_{l}|W_{j}\geq t^{(l)}\}]}\cdot\frac{V^{(l)}_{\mathcal{H}_{0},+}(t^{(l)})}{1+V^{(l)}_{\mathcal{H}_{0},-}(t^{(l)})}\Biggr{\}}$
$\leq$ $\mathbf{E}\Biggl{\{}\sum_{l=1}^{\infty}\frac{v_{l}\alpha}{1.93}\cdot\frac{V^{(l)}_{\mathcal{H}_{0},+}(t^{(l)})}{1+V^{(l)}_{\mathcal{H}_{0},-}(t^{(l)})}\Biggr{\}}$
$\leq$ $\mathbf{E}\Biggl{\{}\sum_{l=1}^{\infty}\frac{v_{l}\alpha}{1.93}\cdot\sup_{t>0}\frac{V^{(l)}_{\mathcal{H}_{0},+}(t)}{1+V^{(l)}_{\mathcal{H}_{0},-}(t)}\Biggr{\}}$
$=$ $\sum_{l=1}^{\infty}\frac{v_{l}\alpha}{1.93}\cdot\mathbf{E}\Biggl{\{}\sup_{t>0}\frac{V^{(l)}_{\mathcal{H}_{0},+}(t)}{1+V^{(l)}_{\mathcal{H}_{0},-}(t)}\Biggr{\}}$
$=$ $\sum_{l=1}^{\infty}{v_{l}\alpha}$
$=$ $\alpha,$ (21)
by (20).
Appendix C Extensions
C.1 Multiple Knockoffs
Although Algorithm 2 with group knockoffs can obtain the rejection set with FDR control, randomness in group knockoffs generation could produce greatly different $\mathcal{R}^{(\text{fg})}$ in different runs, especially in the case that the number of false $H_{j}^{(\text{fg})}$’s is small and Algorithm 2 could return an empty rejection set (Gimenez and Zou,, 2019). Here we extend the proposed Algorithm 2 to multiple group knockoffs $\widetilde{\textbf{X}}^{(1)},\ldots,\widetilde{\textbf{X}}^{(M)}$ where original and knockoff features are simultaneously exchangeable at the level of feature groups. That is to say, with the convention that $\textbf{X}=\widetilde{\textbf{X}}^{(0)}$, $\widetilde{\textbf{X}}^{(0)},\widetilde{\textbf{X}}^{(1)},\ldots,\widetilde{\textbf{X}}^{(M)}$ satisfy
$(\widetilde{\textbf{X}}^{(\sigma_{1}(0))}_{B_{1}},\ldots,\widetilde{\textbf{X}}^{(\sigma_{K}(0))}_{B_{K}},\widetilde{\textbf{X}}^{(\sigma_{1}(1))}_{B_{1}},\ldots,\widetilde{\textbf{X}}^{(\sigma_{K}(1))}_{B_{K}},\ldots,\widetilde{\textbf{X}}^{(\sigma_{1}(M))}_{B_{1}},\ldots,\widetilde{\textbf{X}}^{(\sigma_{K}(M))}_{B_{K}})$ (22)
[…]
for any permutations $\sigma_{1},\ldots,\sigma_{K}$ of $\{0,1,\ldots,M\}$. With importance scores $\{T_{j}^{(m)}|j=1,\ldots,p;m=0,\ldots,M\}$ obtained in an analogous way of (5), we follow He et al., (2021) to compute feature statistics
$\kappa_{j}=\arg\max_{m}T_{j}^{(m)},\quad\tau_{j}=\max_{m}T_{j}^{(m)}-\text{\rm median}\{T_{j}^{(m)}|m\neq\kappa_{j}\},\quad j=1,\ldots,p.$ (23)
Specifically, $\kappa_{j}$ and $\tau_{j}$ are the multiple group knockoffs counterparts of $\text{sign}(W_{j})$ and $|W_{j}|$ respectively (Gimenez and Zou,, 2019) that
  1. $\star$
    conditional on $\tau_{1},\ldots,\tau_{p}$ and $\kappa_{j}$’s for those false $H^{(\text{fg})}_{j}$’s,
    1. $\diamond$
      (Uniformity) $\kappa_{j}$ uniformly distributes on $\{0,1,\ldots,M\}$ for all $j\in\mathcal{H}_{0}^{(\text{fg})}$;
    2. $\diamond$
      (Between-Group Independence) for any $k\neq k^{\dagger}$, $\kappa_{j}$ and $\kappa_{j^{\dagger}}$ are independent for any $j\in B_{k}\cap\mathcal{H}_{0}^{(\text{fg})}$ and $j^{\dagger}\in B_{k^{\dagger}}\cap\mathcal{H}_{0}^{(\text{fg})}$.
Based on $\kappa_{j}$’s and $\tau_{j}$’s, we provide Algorithm 3 as the multiple knockoffs counterpart of Algorithm 2.
Algorithm 3 Feature filter with multiple group knockoffs (naive version).
1:  Input: Groups $B_{1},\ldots,B_{K}$, feature statistics $\{(\kappa_{j},\tau_{j})|j=1,\ldots,p\}$ and the target level $\alpha>0$.
2:  Align $(\kappa_{j},\tau_{j})$’s as shown in Table 6 such that $(\kappa_{j},\tau_{j})$’s for all features in the group $B_{k}$ are in the $k$-th column and $\tau_{(k1)}\geq\tau_{(k2)}\geq\cdots$.
3:  Compute budget $v_{l}$ for the $l$-th row using $\tau_{1},\ldots,\tau_{p}$ for $l=1,2,\ldots$ such that $\sum_{l=1}^{\infty}v_{l}=1$.
4:  Compute grids $\mathcal{G}_{l}=\{1/v_{l},2/v_{l},\ldots,({neg}_{l}+1)/v_{l}\}$ for the $l$-th row where $neg_{l}=\#\{j\in\mathcal{C}_{l}|\kappa_{j}\neq 0\}$.
5:  Combine $\mathcal{G}_{comb}=(\cup_{l}\mathcal{G}_{l})\cup\{0\}$ and sort values ${grid}_{(1)}>{grid}_{(2)}>\ldots$ in $\mathcal{G}_{comb}$.
6:  Initialize $b=0$
7:  repeat
8:     Update $b=b+1$.
9:     Compute
$t^{(l)}=\min\Biggl{\{}t>0\Bigg{|}\frac{1+\#\{j\in\mathcal{C}_{l}|\kappa_{j}\neq 0,\tau_{j}\geq t\}}{v_{l}}\leq{grid}_{(b)}\Biggr{\}},$ (24)
for $l=1,2,\ldots$.
10:  until $\frac{\text{I}\left(\max_{j\in\mathcal{C}_{l}}|W_{j}|\geq t\right)}{M}\times\frac{1+\#\{j\in\mathcal{C}_{l}|\kappa_{j}\neq 0,\tau_{j}\geq t^{(l)}\}}{1\vee[\sum_{l}\#\{j\in\mathcal{C}_{l}|\kappa_{j}=0,\tau_{j}\geq t^{(l)}\}]}\leq\frac{v_{l}\alpha}{1.93}$ for all $l$.
11:  Output: The rejection set $\mathcal{R}^{(\text{fg})}=\cup_{l}\{j\in\mathcal{C}_{l}|\kappa_{j}=0,\tau_{j}\geq t^{(l)}\}$.
Table 6: Alignment of feature statistics $(\kappa_{j},\tau_{j})$’s such that different groups correspond to different columns. Here, $(\kappa_{(kl)},\tau_{(kl)})$ corresponds to the feature from the $k$-th group that is aligned in the $l$-th row.
Row $B_{1}$ $B_{2}$ $B_{3}$cdots
1 $(\kappa_{(11)},\tau_{(11)})$ $(\kappa_{(21)},\tau_{(21)})$ $(\kappa_{(31)},\tau_{(31)})$cdots
[…]
C.2 Filters based on E-values
Another variation relies on the connection between knockoff filter and $e$-values discussed in Ren and Barber, (2024). Following Ren and Barber, (2024)’s way of computing $e$-values of a set of hypotheses whose $W_{j}$’s possess the coin-flipping property, we can analogously compute $e$-values of hypotheses whose $W_{j}$’s are aligned in the same row of Table 2 as shown in steps 4-7 of Algorithm 4. By doing so, an e-value $e_{j}$ is calculated for each hypothesis $H_{j}^{(\text{fg})}$ and we obtain the rejection set $\mathcal{R}^{(\text{fg})}$ by applying the e-BH procedure (Wang and Ramdas,, 2022) to $e$-values $e_{1},\ldots,e_{p}$ (as shown in steps 8-10 of Algorithm 4).
Algorithm 4 Rigorous feature filter with group knockoffs using e-values.
1:  Input: Groups $B_{1},\ldots,B_{K}$, feature statistics $W_{1},\ldots,W_{p}$ and the target level $\alpha>0$.
2:  Align $W_{j}$’s as shown in Table 2 such that $W_{j}$’s for all features in the group $B_{k}$ are in the $k$-th column and $|W_{(k1)}|\geq|W_{(k2)}|\geq\cdots$.
3:  Compute the number of feature $p_{l}$ in the $l$-th row ($l=1,2,\ldots$).
4:  for $l=1,2,\ldots$ do
5:     Compute
$t_{l,\alpha}=\min\Biggl{\{}t_{l}>0\Bigg{|}\frac{1+\#\{j\in\mathcal{C}_{l}|W_{j}\leq-t_{l}\}}{1\vee\#\{j\in\mathcal{C}_{l}|W_{j}\geq t_{l}\}}\leq\alpha/2\Biggr{\}}.$
6:     Compute e-value $e_{j}$ for each $j\in\mathcal{C}_{l}$ as
$e_{j}=\frac{p_{l}\times I(W_{j}\geq t_{l,\alpha})}{1+\#\{j\in\mathcal{C}_{l}|W_{j}\leq-t_{l,\alpha}\}}.$
7:  end for
8:  Compute $e_{(1)}\geq\ldots\geq e_{(p)}$ as the order statistics of $e_{1},\ldots,e_{p}$.
9:  Compute $\widehat{J}=\max\{J|e_{(J)}\geq p/(\alpha J)\}$ or $\widehat{J}=0$ if $e_{(J)}<p/(\alpha J)$ for $J=1,\ldots,p$.
10:  Output: The rejection set $\mathcal{R}^{(\text{fg})}=\{j|e_{j}\geq e_{(\widehat{J})}\}$.
Theorem 3.
The rejection set $\mathcal{R}^{(\text{fg})}$ obtained by Algorithm 4 controls $\text{FDR}^{(\text{fg})}$ at the target level $\alpha>0$.
Proof of Theorem 3.
Because $W_{j}$’s that are in the same row of Table 2 possess the coin-flipping property, by (7) of Ren and Barber, (2024), we have
$\textbf{E}\left\{\frac{\sum_{j\in\mathcal{C}_{l}\cap\mathcal{H}_{0}^{(\text{fg})}}I(W_{j}\geq t_{l,\alpha})}{1+\sum_{j\in\mathcal{C}_{l}\cap\mathcal{H}_{0}^{(\text{fg})}}I(W_{j}\leq-t_{l,\alpha})}\right\}\leq 1.$
As a result, we have for $l=1,2,\ldots,$
$\sum_{j\in\mathcal{C}_{l}\cap\mathcal{H}_{0}^{(\text{fg})}}\textbf{E}\left\{e_{j}\right\}=$ $p_{l}\times\textbf{E}\left\{\frac{\sum_{j\in\mathcal{C}_{l}\cap\mathcal{H}_{0}^{(\text{fg})}}I(W_{j}\geq t_{l,\alpha})}{1+\sum_{j\in\mathcal{C}_{l}}I(W_{j}\leq-t_{l,\alpha})}\right\}$
$\leq$ $p_{l}\times\textbf{E}\left\{\frac{\sum_{j\in\mathcal{C}_{l}\cap\mathcal{H}_{0}^{(\text{fg})}}I(W_{j}\geq t_{l,\alpha})}{1+\sum_{j\in\mathcal{C}_{l}\cap\mathcal{H}_{0}^{(\text{fg})}}I(W_{j}\leq-t_{l,\alpha})}\right\}$
$\leq$ $p_{l},$
and thus
$\sum_{j\in\mathcal{H}_{0}^{(\text{fg})}}\textbf{E}\left\{e_{j}\right\}=\sum_{l=1}^{\infty}\sum_{j\in\mathcal{C}_{l}\cap\mathcal{H}_{0}^{(\text{fg})}}\textbf{E}\left\{e_{j}\right\}\leq\sum_{l=1}^{\infty}p_{l}=p.$
Finally, by Theorem 2 of of Ren and Barber, (2024), as $\sum_{j\in\mathcal{H}_{0}^{(\text{fg})}}\textbf{E}\left\{e_{j}\right\}\leq p$, the e-BH procedure in steps 8-10 of Algorithm 4 returns the rejection set $\mathcal{R}^{(\text{fg})}$ with FDR control. ∎